Tech Staff Meeting Notes December 1, 2015

The top priority is always repair of down stations.
NM problem stations
ET problem stations

Duty Table updated June 2014


News & Highlights

The CERI Holiday party will be on Friday December 11. Historically its been around lunch time. Presumably it will be pot luck. There should be an announcement forthcoming from Michelle.

James was pleased to announce that the decommission of folkworm is complete. (and he did it in a fairly short amount of time). Mitch wants the computer for a souvenir in his office. We're still dependent on the old EIDS/QDM feed running on shakeworm. Mitch will send email to James and Steve to investigate replacing that possibly with a geojson feed.


Put AO/TA conversions on hold for now. Jim requested the extra sim card now so that it could be used for several noise studies on deck. The six stations in the AO network are being decommissioned and the hardware returned to the ANSS Depot. CERI offered to take over O&M for two: U40A (near Mountain Home, AR) and X40A (east of Hot Springs). Jim did a recon on U40A and its doable but will take a little work. This effort will likely take about 6 months.

2015 Preventive Maintenance84% complete
NM BB/SM & SP: 67 done & 10 to go
NM SM: 9 done & 8 to go
AG BB/SM: 6 done & 0 to go
GPS: 13 done & 0 to go
Misc: 6 done & 2 to go

freewave serial reset installed at 10 of 11 broadbands. GLAT remains.


Data taken from:

NM Down

AG Down
FCAR.HH* -sporadic
FCAR.HN* - "
HHAR.HH* - "
HHAR.HN* - "

NM Misc Problems
CPAR - spikes
HCAR - "
TYAR - "
HDBT.HNE - see image

MSAR has slumped down the hill and the pipe vault is now bent. Need to install a repeater to get the outer station, RVAR, back online until access improves enough to rebuild MSAR.

All ISIS3 hardware except that slated for PLAR is installed. Have about 26 remaining, mostly MKTA and some NMAD outer stations.

RDST has a very large pole right next to the vault that will corrupt any data during strong ground motion. It needs to get moved; Dave has a couple options on the Rhodes campus (the name will change when its moved). We have a Basalt almost ready to go so we and install new rather than move old.

NODE Tasks

The win7 upgrades at nodes (mudworm is done). There are several win7 issues that Steve and James are working out before installing at remote locations.


The Wilder-SFTN DBB link will be converted from aironets to the 2.5 GHz Ubiquiti Bullet M radios.
They will do 600 mW which should give us +6 dBm of signal strength at both ends which would negate the need to go from a 3' dish at Shelby to a 6 ft. That link is the bottleneck and these radios are rated at 100 Mbps so they have way more bandwidth than we need. They play well with the aironets and the conversion at Wilder's and Shelby would be simple. They screw onto the bulkhead mounted TSP in the existing enclosure. The pair of Bullet's cost $160. This will get done Wednesday 12/2.

There are some inconsistencies in the configuratons of the aironets. Steve will review and sync configs.

The Luxora link was getting considerable interference causing degraded signal and connection loss. Jim changed frequencies and it got better. Consider this as a candidate for a different radio.




The waveserver continues to have occasional very gappy periods. James found that rtpm was configured to store sms messages and there were an extremely large number of those files. They were cleaned up and reconfigured on all nodes. So far nmad seems to have improved and no significant gaps were produced over the weekend.

AZTN/HDBR: Shahram submitted a proposal to TDOT to replace all the sensors with Episensors (including new cabinets, etc). We asked USGS to provide new radios so we can have a continuous buffer of all channels at AZTN. We'll then allow access to NSMP so they can extract event data. Unofficial word is TDOT approved the funding; waiting on signed contract.


Need to review and update the power systems in 218. Meeting with IT staff is scheduled for 12/3.


Replacement station for GTTN and RBNC being evaluated and permissioned for Southern Highlands Science Learning Center at Purchase Knob, a part of the Great Smoky Mountains National Park. Need to do a noise study.


Taken from:

ET Down

Misc Problems
CPRT.HNN - drifting
GMG - connection issues?

John replaced the lengthy list below with a prioritized spreadsheet HERE .


Improve RCGA AC power and perform maintenance work on shed.

Replace BCRT Strong Motion Vault and instrument. Improve transient suppression/g rounding.

Replace GMG RT-147.

Move PMTN due to excessive noise and vandalism. The trees at Lone Mountian are t oo tall to get a signal out. Looking for an alternate site. Can we use a cell modem?


HPNC node is up. Need to earthworm commands (e.g. status, restart etc) from the ssh command line; currently remote desktop is required to interface with earthworm. Need to try to get 3 complete 3-c stations operating: MGNC, GFM and SMNC. Export those channels along with picks and migrate Mt Gibbs node (BRBC) helicorder to HPNC. Done, SMNC is now DSNC. GFM, MGNC, and DSNC are updating in heli.

Upgrade remaining High Peak Node Stations from L4-3D's to S-13's. At least one bad component in L4-3's at CCNC, TVNC, TRYN.

GFM needs to be upgraded to ISIS4. Done. Need to do a noise study for a closer site, then move it in the spring along with upgrading to S-13.


Replace VHTN E-W and N-S Instruments.

WMTN pit needs to be moved due to erosion.

CCNC is not going through fgtn and instead repeats through ASTN. The 950MHz shot would work so it could be converted to digital.

We're still repeating CPRT through GTTN. The new internet host at the Corryton Senior Center for CPRT has been verbally approved by Greg Hill the center coordinator and the Knox County facility supervisor.


No problems at this time.

AG Network

HHAR needs a new router and power supply. Move the solar to a location with more sun.

FCAR AC power upgrade was abandoned. A tree was removed improving sunlight but still needs additional solar.


Needs some clean up.

The FDL wiki is here:


OXF and LRAL are still down. Depot personnel were scheduled to make sites visits but haven't yet. Mitch will check and offer assistance.

Giddings repairs and train Randy's student prior to 12/11-1/15 deployment.

The seisnet group on umdrive is at You will need your uuid and password.

The seisnet wiki is at You will need your uuid and password.